OpenStudy (anonymous):

Passing an electric current through a sample of water (H2O) can cause the water to decompose into hydrogen gas (H2) and oxygen gas (O2) according to the following equation. 2H2O= 2H2 + O2 The molar mass of H2O is 18.01 g/mol. The molar mass of O2 is 32.00 g/mol. What mass of H2O, in grams, must react to produce 50.00 g of O2? 14.07 23.05 28.14 56.28

OpenStudy (anonymous):

@Abmon98

OpenStudy (anonymous):

is it 23.05

OpenStudy (abmon98):

nope your given a know mass of oxygen which is 50 grams Number of moles=mas/molar mass 50/32=1.5625 moles of oxygen look for the mole ratio of H2O to O2 which is 2:1 1.5625*2/1=3.125 moles of H2O 3.125*18=Mass

OpenStudy (anonymous):

so 56.28 ?

OpenStudy (anonymous):

is that my answer